# Who Can Withstand Chat-Audio Attacks? An Evaluation Benchmark for Large Language Models

### v1

@@ 6 Conclusion
This work explored the vulnerabilities of large language models (LLMs) to adversarial audio attacks in conversational scenarios. We introduced the Chat-Audio Attacks (CAA) benchmark, consisting of 360 adversarial attack sets across four attack types: content, emotional, explicit noise, and implicit noise attacks. Our evaluation of six state-of-the-art LLMs using three methods—Standard Evaluation, GPT-4o-Based Evaluation, and Human Evaluation—revealed and discussed significant model vulnerabilities under adversarial conditions.
The CAA benchmark highlights these weaknesses and provides a foundation for developing more robust defense mechanisms. As LLMs are increasingly integrated into voice interactions, enhancing their resilience against adversarial audio attacks remains a crucial area for future research.
@@ 7 Limitations
Meanwhile, there is very limited research on audio jailbreak attacks, and there is almost no open-source work available in this area. As a result, we were unable to generate corresponding adversarial samples for testing such attacks. This represents a significant gap in the exploration of both white-box and black-box attacks on specific large language models (LLMs). There is considerable room for future development in addressing the security vulnerabilities posed by these types of attacks, which remain an important but underexplored aspect of LLM security.
@@ 8 Ethics Statement
This research explores vulnerabilities in large language models (LLMs) through adversarial audio attacks with the goal of improving model robustness. All adversarial samples were generated solely for research purposes to enhance the security of LLM-based systems. Human evaluation was conducted with informed consent, and no personally identifiable information was collected. We prioritize ethical considerations, ensuring that the work contributes to safer and more reliable conversational AI technologies.
@@ Appendix C Qualitative Results
Table 6 provides examples of responses generated by the six multimodal audio LLMs when faced with different adversarial samples. It illustrates the varying impacts of adversarial attacks on each model, clearly highlighting the degree to which different models are affected.
@@
Table 6: Examples of response generated by LLMs. Blue indicates abnormal response.

### v2

@@ 6 Conclusion
This work explored the vulnerabilities of LALMs to adversarial audio attacks in conversational scenarios. We introduced the Chat-Audio Attacks (CAA) benchmark, consisting of 360 adversarial attack sets across four attack types: content, emotional, explicit noise, and implicit noise attacks. Our evaluation of six state-of-the-art LALMs using three methods—Standard Evaluation, GPT-4o-Based Evaluation, and Human Evaluation—revealed and discussed significant model vulnerabilities under adversarial conditions.
The CAA benchmark highlights these weaknesses and provides a foundation for developing more robust defense mechanisms. As LALMs are increasingly integrated into voice interactions, enhancing their resilience against adversarial audio attacks remains a crucial area for future research.
@@ 7 Limitations
Meanwhile, there is very limited research on audio jailbreak attacks, and there is almost no open-source work available in this area. As a result, we were unable to generate corresponding adversarial samples for testing such attacks. This represents a significant gap in the exploration of both white-box and black-box attacks on specific LALMs. There is considerable room for future development in addressing the security vulnerabilities posed by these types of attacks, which remain an important but underexplored aspect of LALM security.
@@ 8 Ethics Statement
This research explores vulnerabilities in LALMs through adversarial audio attacks with the goal of improving model robustness. All adversarial samples were generated solely for research purposes to enhance the security of LALM-based systems. Human evaluation was conducted with informed consent, and no personally identifiable information was collected. We prioritize ethical considerations, ensuring that the work contributes to safer and more reliable conversational AI technologies.
9 Acknowledgements
This project is partially supported by ARC DP240101349.
@@ Appendix C Qualitative Results
Table 6 provides examples of responses generated by the six LALMs when faced with different adversarial samples. It illustrates the varying impacts of adversarial attacks on each model, clearly highlighting the degree to which different models are affected.
@@
Table 6: Examples of response generated by LALMs. Blue indicates abnormal response.